OpenStudy (anonymous):

Solve. 2x^2 + x – 1 = 2

OpenStudy (anonymous):

Choices: http://awesomescreenshot.com/0c71tknh96

OpenStudy (anonymous):

Is 2x2 multiplying as 2*2?

OpenStudy (anonymous):

x=-1 If not

OpenStudy (e.cociuba):

No, I believe it is 2x^2.

OpenStudy (anonymous):

Yes @e.cociuba

OpenStudy (anonymous):

{x=(-3/2), x=1}

OpenStudy (anonymous):

I showed you the quadratic formula already. \[x = \frac{-b \pm \sqrt{b^{2} - 4ac}}{2a}\] You can solve any quadratic equation with that!

OpenStudy (e.cociuba):

@LaceyLeanne explain it. Don't just give the answers you found on the internet. He needs to be helped and guided into solving the problem.

OpenStudy (anonymous):

FIRST MISTAKE: Found on the internet SECOND MISTAKE: He has a formula, he can figure it out!